# Model Predictive Control via Probabilistic Inference: A Tutorial

### v1

@@ 3.2 Analysis of the Optimal Control Distribution $\pi^{*}$
Having derived the optimal control distribution in Eq. (2) through mathematical manipulations, we now pause to examine its meaning and characteristics.
@@ 3.2.1 Influence of the Prior Distribution on the Optimal Control Distribution
As shown in Eq. (2), the optimal control distribution is expressed as the product of the Boltzmann distribution and the control input prior distribution. This means that “optimal actions” are not merely those minimizing the cost function in Eq. (1a), but rather those that also account for the action prior distribution (Fig. 5).
Reference [kappen2005path] illustrates this characteristic with the analogy of a “drunken spider crossing a pond.” Imagine a spider attempting to cross a pond via a narrow bridge. When sober (i.e., when the prior distribution has small variance), the optimal action is to cross the pond via the shortest path using the bridge. However, when drunk (i.e., when the prior distribution has large variance), there is a risk of falling into the pond while crossing the narrow bridge. Therefore, detouring around the pond becomes the optimal action. Thus, optimal actions depend not only on the cost function but also on the prior distribution, which appears to be a principle valid even in nature.
Note that models that use Boltzmann distributions as optimal distributions for sample generation are called Energy-Based Models (EBM), long studied in machine learning [lecun2006tutorial]. The difference between EBM and Eq. (2) lies in the prior distribution, which originates from the dynamics in Eq. (1b).
@@ 3.2.2 Decision-Making Through Symmetry Breaking
Figure 6: Example of decision-making through symmetry breaking.
@@
Consider the obstacle avoidance example illustrated in Fig. 6. When far from the obstacle, the optimal control distribution has two modes representing paths around the obstacle. At this point, the probabilities of selecting left or right paths are nearly equal, resulting in no execution of avoidance behavior. However, as time progresses and the robot approaches the obstacle, at some point the prior distribution randomly biases toward one direction. Subsequently, the optimal control distribution, bound to the prior, converges to a single mode over time.
@@ 3.2.3 Relationship Between Temperature Parameter and Optimal Control Distribution
Figure 5 illustrates how the optimal control distribution changes with the temperature parameter. Larger temperature parameters increase the influence of the prior distribution, making the optimal control distribution smoother and closer to the prior. Conversely, smaller temperature parameters increase the influence of the Boltzmann distribution, producing sharper distributions concentrated near the Boltzmann distribution’s modes. As extreme cases, $\lambda\to 0$ yields deterministic optimal actions that minimize the cost function, while $\lambda\to\infty$ yields stochastic optimal actions following only the prior distribution. Thus, the temperature parameter tunes the balance between deterministic and stochastic characteristics of the optimal control distribution.
@@ 3.2.4 Factors Governing Sample Complexity
As we will discuss in Section 3.3, we ultimately approximate the optimal control distribution in Eq. (2) using Monte Carlo methods. Therefore, the optimal control distribution should ideally have a shape that can be approximated with as few samples as possible. The number of samples required to approximate a probability distribution depends on its shape, a property known as sample complexity.
Generally, sample complexity increases as the distribution becomes more peaked, requiring more samples [^4] Intuitively, drawing from a broader “winning region” with random samples requires fewer samples than pinpointing a single “correct answer.”. According to Eq. (2), the main factors determining the optimal control distribution’s shape are:
  • Temperature parameter $\lambda$: As discussed in Section 3.2.3, smaller temperature parameters produce sharper optimal control distributions with higher sample complexity.
  • Prior distribution $p(\mathbf{u}_{0:T-1})$: Sharper priors (i.e., smaller exploration noise) increase sample complexity.
  • Cost function $J_{\tau}$: When the cost function landscape is steep, e.g., when many obstacles lie directly in front of the robot such that many samples collide, sample complexity increases.
Since these interacting factors make estimating required number of samples difficult, in practice, more samples are generally better.
@@ 3.3 Control Input Generation from the Optimal Control Distribution: The MPPI Example
Despite having a mathematical expression for the optimal control distribution in Eq. (2), directly sampling from it computationally is challenging. This is because the distribution can be highly complex depending on the cost function and prior, or may be ill-defined until a specific control input sequence is determined. We therefore again employ a variational inference approach. Specifically, we approximate the optimal control distribution $\pi^{*}$ using a simple variational distribution $\pi_{\theta}$:
@@
and find the variational distribution $\pi_{\theta}$ that achieves this minimum. The specific choice of $\pi_{\theta}$ varies by method.
The most widely used probabilistic inference MPC method is MPPI [williams2018information]. MPPI employs a multivariate Gaussian distribution with fixed covariance $\Sigma$ as the variational distribution. Specifically, setting $\theta=\mu_{0:T-1}=[\mu_{0},\dots,\mu_{T-1}]$ and
@@
Since Eq. (10) is concave, the global optimal solution $\mu_{0:T-1}^{*}$ maximizing Eq. (11) can be found via the stationarity condition. Setting $\mathcal{L}=\int\pi^{*}\log\mathcal{N}(\mu_{t},\Sigma)d\mathbf{u}_{0:T-1}$:
@@
    Sample $K$ control inputs from the prior distribution:
@@
    Forward simulation and calculating trajectory costs using Eq. (7):
@@
    Find an optimal solution via Monte Carlo approximation of Eq. (12):
@@
Note that Eq. (12) can be implemented simply using the Softmax function as shown in Eq. (13), i.e., merely a weighted average of input samples using Softmax weights. Moreover, steps (1) and (2) are independent operations for each sample $k$, enabling straightforward GPU parallelization via PyTorch and similar frameworks [kohonda2025mppi_playground].
A key advantage of MPPI is that it finds the global optimal solution (in the sense of KL divergence minimization) through a single round of sampling and weighted averaging [^5] However, as discussed in Section 3.2.4, the required number of samples depends on the configuration.. This stems from the concavity of Eq. (10), which results from fixing the variance of the multivariate Gaussian variational distribution. If we simultaneously optimize the variance, Eq. (11) would no longer be concave, requiring iterative repetition of the three steps above in a single optimization, increasing computational cost. However, since the variational distribution’s variance setting significantly impacts performance, careful configuration is necessary [honda2024stein].
4 Applications and Extensions
The previous sections presented the theoretical framework of probabilistic inference-based MPC. The general scheme can be summarized by the following equation:

### v2

@@ 2.4 Analysis of the Optimal Control Distribution $\pi^{*}$
Having derived Eq. (2), we now briefly discuss its interpretation and key properties.
@@ 2.4.1 Influence of the Prior Distribution on the Optimal Control Distribution
As shown in Eq. (2), the optimal control distribution is expressed as the product of the Boltzmann distribution and the control input prior distribution. Thus, “optimal actions” are not simply those that minimize the cost in Eq. (1a); they also reflect the action prior (Fig. 5).
Reference [kappen2005path] illustrates this characteristic with the analogy of a “drunken spider crossing a pond.” Imagine a spider attempting to cross a pond via a narrow bridge. When sober (i.e., when the prior has small variance), the best strategy is to take the shortest route across the bridge. When drunk (i.e., when the prior variance is large), crossing the narrow bridge becomes risky, and a longer detour around the pond may be preferable. The lesson is that optimal actions depend not only on the cost function but also on the prior distribution.
Note that models that use Boltzmann distributions as optimal distributions for sample generation are called Energy-Based Models (EBM), well studied in machine learning [lecun2006tutorial]. Compared with standard EBMs, Eq. (2) includes an explicit action prior, which is induced by the system dynamics in Eq. (1b).
@@ 2.4.2 Decision-Making Through Symmetry Breaking
Figure 6: Decision-making through symmetry breaking in the optimal control distribution. When multiple avoidance paths exist, the optimal control distribution initially exhibits multiple modes with nearly equal probability. As time progresses and the prior variance effectively decreases, stochastic bias toward one direction causes the distribution to collapse into a single mode, resulting in delayed but decisive action selection.
@@
Consider the obstacle-avoidance example in Fig. 6. When the robot is far from the obstacle, the optimal control distribution is typically bimodal, corresponding to left and right avoidance paths. The two modes often have nearly equal probability, so the controller does not commit strongly to either direction. As the robot approaches the obstacle, random fluctuations in the prior can bias the distribution toward one side, and the optimal control distribution (constrained by the prior) eventually collapses to a single mode.
@@ 2.4.3 Relationship Between Temperature Parameter and Optimal Control Distribution
Figure 5 illustrates the effect of $\lambda$. Larger values of $\lambda$ increase the influence of the prior, yielding smoother distributions closer to the prior. Smaller values emphasize the Boltzmann term, producing sharper distributions concentrated near low-cost regions. As extreme cases, $\lambda\to 0$ yields deterministic optimal actions that minimize the cost function, while $\lambda\to\infty$ yields stochastic optimal actions following only the prior distribution. Thus, the temperature parameter tunes the balance between deterministic and stochastic characteristics of the optimal control distribution.
The temperature parameter is a hyperparameter that must be tuned for each application. Smaller $\lambda$ encourages more aggressively cost-seeking behavior, which can improve performance but may reduce sample efficiency and robustness to disturbances. Larger $\lambda$ places more weight on the prior, which can improve sample efficiency and robustness but may become overly conservative. In many works, a fixed temperature parameter is manually tuned, though adaptive adjustment methods have also been proposed [watson2023inferring, pezzato2025sampling].
@@ 2.4.4 Factors Governing Sample Complexity
As we discuss in Section 2.5, in practice we approximate the optimal control distribution in Eq. (2) using Monte Carlo sampling. Accordingly, it is desirable for the target distribution to be well approximated with as few samples as possible. The number of samples needed to obtain a good approximation depends strongly on the distribution’s shape; this is often referred to as sample complexity.
In general, sample complexity increases as the distribution becomes more sharply peaked, requiring more samples [^5] Intuitively, it is easier to hit a broad “winning region” with random samples than to pinpoint a single “correct answer.”. According to Eq. (2), the main factors determining the optimal control distribution’s shape are:
  • Temperature parameter $\lambda$: as discussed in Section 2.4.3, smaller $\lambda$ yields sharper distributions and thus higher sample complexity.
  • Prior distribution $p(\mathbf{u}_{0:T-1})$: tighter priors (i.e., less exploration noise) also increase sample complexity.
  • Cost landscape $J_{\tau}$: when the cost changes steeply (e.g., many samples collide with obstacles), the effective distribution becomes harder to approximate, increasing sample complexity.
Because these factors interact, it is difficult to predict the required number of samples a priori; in practice, using more samples is often beneficial when computationally feasible.
@@ 2.5 Control Input Generation from the Optimal Control Distribution: The MPPI Example
Despite having a mathematical expression for the optimal control distribution in Eq. (2), directly sampling from it computationally is challenging. This is because the distribution can be highly complex depending on the cost function and prior, or may be ill-defined until a specific control input sequence is determined. We therefore again employ a VI approach. Specifically, we approximate the optimal control distribution $\pi^{*}$ using a parameterized variational distribution $\pi_{\theta}$:
@@
and seek the variational distribution $\pi_{\theta}$ that achieves this minimum. The specific choice of $\pi_{\theta}$ varies by method, as discussed in Section 3.2.1.
The most widely used PI-MPC method is MPPI [williams2018information]. MPPI employs a multivariate Gaussian distribution with fixed covariance $\Sigma$ as the variational distribution. Specifically, setting $\theta=\mu_{0:T-1}=[\mu_{0},\dots,\mu_{T-1}]$ and
@@
Since Eq. (10) is concave, the global optimal solution $\mu_{0:T-1}^{*}$ can be obtained via the stationarity condition. Setting $\mathcal{L}=\int\pi^{*}\log\mathcal{N}(\mu_{t},\Sigma)d\mathbf{u}_{0:T-1}$:
@@
    Sample $K$ control input sequences from the prior distribution:
@@
    Forward rollout and trajecotry cost evaluation using Eq. (7):
@@
    Update the mean via the Monte Carlo approximation of Eq. (12):
@@
Note that Eq. (12) can be implemented using the Softmax weighting in Eq. (13), i.e., a weighted average of the sampled control sequences. Moreover, steps (1) and (2) are independent operations for each sample $k$, enabling straightforward GPU parallelization via PyTorch and similar frameworks. If input constraints must be enforced, we can clamp each sampled control sequence between steps (1) and (2); this clamping can be treated as part of the dynamics.
A key advantage of MPPI is that it finds the global optimal solution (in the sense of KL divergence minimization) through a single round of sampling and weighted averaging [^6] However, as discussed in Section 2.4.4, the required number of samples depends on the configuration.. This stems from the concavity of Eq. (10), which results from fixing the covariance of the multivariate Gaussian variational distribution. If we simultaneously optimize the covariance, Eq. (11) would no longer be concave, requiring iterative repetition of the three steps above in a single optimization, increasing computational cost. However, since the variational distribution’s covariance setting significantly impacts performance, careful configuration is necessary, and discussed in Section 3.1.
2.6 Summary of the Tutorial
In this tutorial, we introduced the basic theory behind PI-MPC. We conclude by summarizing the framework with the following relation: